OpenStudy (anonymous):

can someone tell me what im doing wrong? When I enter this into the calculator I get 4.08, but the site I'm using to study has 3.126 as an answer. 2x20xsin50^0/9.8 = 3.126 s.

OpenStudy (ahsome):

Equation is: \(2*20*sin50^0/9.8\) Is this it?

OpenStudy (anonymous):

yeah @Ahsome

OpenStudy (kropot72):

My calculator gives the result as 3.126712

OpenStudy (kropot72):

I calculated the following: \[\large \frac{2\times20\times \sin50^{0}}{9.8}\]

OpenStudy (anonymous):

this time i did not add a exponent to my sin(50) and I got 3.126. Am I supposed to not manually do sin(50)^0 in the calculator or something?

OpenStudy (kropot72):

The expression \[\large \sin 50^{0}=\sin50\ degrees\] The symbol following sin 50 is not an exponent, it is an accepted symbol for 'degrees'

OpenStudy (kropot72):

You now are getting the correct result, therefore you must have set your calculator to the 'degrees' mode. When you enter 'sin 50 degrees' into your calculator, just press the 'sin' button and then enter '50'. Note: Some calculators need parenthesis added.